# Nested and outlier embeddings into trees

### v1

@@ 2 Definitions and results
Further, we can efficiently sample the from the distribution $\mathcal{D}$ if we can efficiently sample from $\mathcal{D}_{S}$. We will also briefly discuss the implications of this theorem for ultrametric embeddings in Appendix E
Using Theorem 2.4 and a modification of the LP from Munagala et al. [39] we get the following result.
Theorem 2.5:
Given a metric $(X,d)$ and target distortion $c$, let $k$ be the optimal outlier set size for probabilistically embedding $(X,d)$ into HSTs. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
The LP-based techniques we use also allow us to obtain the following result for the weighted variant of the problem, in which we are given weights $w_{i}$ on all nodes $i$ in the metric space. The cost of a subset $K$ of the metric space is $w(K):=\sum_{i\in K}w_{i}$, and the goal is to find an outlier embedding with some target distortion such that the cost of the outlier set is minimized.
Corollary 2.6:
@@
The rest of the paper will proceed as follows:
  • In Section 3, we discuss the buy-at-bulk, dial-a-ride, and minimum communication cost tree problems as algorithmic settings where Theorem 2.5 may be useful. We also discuss the existence of metrics for which removing a small number of points greatly improves their minimum distortion for probabilistic HST embeddings.
  • In Section 4, we show how to generalize the $\ell_{1}$ nested composition algorithm of Chawla and Sheridan [20] and characterize the type of input merge functions that allow this algorithm to work. We also show how the “obliviousness” of this algorithm allows it to work in both deterministic and probabilistic settings.
  • In Section 5, we show how to find a good merge function for HST embeddings, and we show how to modify the algorithm of Munagala et al. [39] such that the existence of good nested compositions for HST embeddings results in a good approximation factor for this modified algorithm for the outlier setting.
Finally, in the appendices we will discuss some of the proofs omitted from the main paper and we will discuss a set of conditions on deterministic Lipschitz extensions that allow them to work in the probabilistic setting.

### v2

@@ 2 Definitions and results
Further, we can efficiently sample from the distribution $\mathcal{D}$ if we can efficiently sample from $\mathcal{D}_{S}$. We will also briefly discuss an extension of this theorem to ultrametrics in Appendix 0.D.
In Section 4, we use Theorem 2.1 and a modification of the LP from Munagala et al. [31] to prove Theorem 1.1. Our LP-based techniques also allow us to obtain the following result for the weighted variant of the problem. [^3] In the weighted outlier set problem we are given weights $w_{i}$ on all nodes $i$ in the metric space. The cost of a subset $K$ of the metric space is $w(K):=\sum_{i\in K}w_{i}$, and the goal is to find an outlier embedding with some target distortion such that the cost of the outlier set is minimized.
Corollary 1
@@
3 Probabilistic nested embeddings
In this section we prove Theorem 2.1 by first establishing a set of properties for the existence of good probabilistic nested embeddings. We then show how to achieve these properties for HST embeddings Recall that we are given a metric space $(X,d_{X})$, a target class of metrics $\mathcal{Y}$, a subset $S$ with $k=|X\setminus S|$, and an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$. Our goal is to produce a single probabilistic embedding for the entire set $X$ that has small expansion.
Our approach is based on the nested embedding technique of Algorithms 4.1 and 5.1 in Chawla and Sheridan [19] for $\ell_{1}$ embeddings but strengthens that technique by introducing a general procedure for merging embeddings. There are two challenges that we need to address in order to adapt these algorithms to our setting. First, we need to ensure that the final composition does not contract distances. In [19]’s Algorithm 4.1, although the expansion for every pair of points can be bounded for this composition, distances can get contracted by arbitrarily large factors – indeed some vertices end up colocated under the composition. As in [19], we randomly partition $K=X\setminus S$ into sets $K_{1},K_{2},\ldots$, and identify a “close” vertex $\gamma_{i}\in S$ for each such set. Then, we consider the embedding in the given assortment that maps $K_{i}\cup\{\gamma_{i}\}$ into $\mathcal{Y}$. We “merge” this embedding with the embedding $\alpha_{S}$ in such a fashion that neither embedding by itself gets distorted any further. This allows us to bound contraction in a similar way to Algorithm 5.1 from [19], though we frame it more generally. We then iterate in this manner over all of the components in the partition of $K$.
The second challenge is to combine the per-pair $\alpha_{S}$ expansion guarantee into a expected distortion guarantee when we draw an $\alpha_{S}$ from the given probabilistic embedding $\mathcal{D}_{S}$. The key property that allows us to do so is the independence of the procedure for partitioning $K$ and the structure of $\alpha_{S}$. We call this property obliviousness and will revisit it more formally in Appendix 0.C.
Definition 4(Perfect merge)
Let $Z_{1},Z_{2}\subseteq X$ with $Z_{1}\cap Z_{2}=\{v\}$. Let $\alpha_{1}:Z_{1}\rightarrow Y_{1}$ and $\alpha_{2}:Z_{2}\rightarrow Y_{2}$ be deterministic expanding embeddings, for $Y_{1},Y_{2}\in\mathcal{Y}$. We say that a procedure Merge is a perfect merge function for embedding $X$ into $\mathcal{Y}$ if given any pair $\alpha_{1}$ and $\alpha_{2}$ of embeddings from subsets of $X$ into $\mathcal{Y}$ satisfying the conditions above, it produces a an embedding $\alpha:Z_{1}\cup Z_{2}\rightarrow Y$ for some $Y\in\mathcal{Y}$ such that:
  1. 1.
    For all $x,y\in Z_{1}$, $d_{\alpha}(x,y)=d_{\alpha_{1}}(x,y)$.
  2. 2.
    For all $x,y\in Z_{2}$, $d_{\alpha}(x,y)=d_{\alpha_{2}}(x,y)$.
  3. 3.
Now we present Algorithm 1 that, given an assortment of probabilistic embeddings on subsets of $X$, produces a new embedding that is drawn from a strongly nested composition of the input embeddings. We highlight in red where Algorithm 1 differs from Algorithm 5.1 in [19]. In [19], the input PerfectMerge function was merely concatenation of vectors in the $\ell_{1}$ metric. We will develop a different PerfectMerge merge function for embeddings into HSTs in this section.
Input: Metric space $(X,d)$, subset $S\subseteq X$, $k=|X\setminus K|$; Assortment of embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ where $\mathcal{D}_{S}:S\rightarrow\mathcal{Y}$ and $\mathcal{D}_{K^{\prime}}:K^{\prime}\rightarrow\mathcal{Y}$ for $K^{\prime}\subset X$; A function PerfectMerge that takes in two deterministic embeddings $\alpha:Z\rightarrow\mathcal{Y}$ and $\alpha^{\prime}:Z^{\prime}\rightarrow\mathcal{Y}$ for $Z,Z^{\prime}\subset X$ and $|Z\cap Z^{\prime}|=1$, and returns an embedding $\alpha^{\prime\prime}:Z\cup Z^{\prime}\rightarrow\mathcal{Y}$.
[…]
2: Sample $\alpha_{S}\leftarrow\mathcal{D}_{S}$
4:Select $b$ uniformly at random from the range $[2,4]$
[…]
7:for $i=1$ to $k$ do
8:   $u_{i}\leftarrow\pi^{-1}(i)$
9:   $K_{i}\leftarrow\{v\in K^{\prime}\ |\ d(v,u_{i})\leq b\cdot d(v,\gamma(v))\}$ $\triangleright$ Let the “center" of $K_{i}$ be $u_{i}$
10:   Sample $\alpha_{i}\leftarrow\mathcal{D}_{K_{i}\cup\set{\gamma(u_{i})}}$
11:   $\alpha\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$
12:  $K^{\prime}\leftarrow K^{\prime}\setminus K_{i}$
13:Output $\alpha$
Theorem 3.1
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 1 returns a sample from a strong probabilistic $(3,g)$-nested composition with $g(c_{S},c_{X},k)=O((c_{S}+c_{X}(k))\cdot H_{k})$, where $H_{k}$ is the $k$th harmonic number and $c_{X}(k)$ is an upper bound on the distortion for any of the $\mathcal{D}_{K^{\prime}}$ in the input.
We break down the proof of the theorem into two parts: bounding the expansion of the embedding (Lemma 1) and its contraction (Lemma 2).
Lemma 1
$\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(x,y)]$ $\leq(125\,c_{S}+116\,c_{X}(k))\cdot H_{k}\cdot d(x,y).$
Further, if $x,y\in S$, then $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(x,y)]\leq c_{S}\cdot d(x,y)$.
Lemma 2
We formally prove each of these lemmas in Appendix 0.A.1 and 0.A.2, respectively. Lemma 1 follows essentially the same proof technique as that of Algorithm 5.1 in [19], though we pay special attention to the independence of the $\alpha_{S}$ we draw from $\mathcal{D}_{S}$ and the random partitioning of $K$. To intuitively understand the contraction bound, note that when some node $v\in K_{i}$ is added to the domain of $\alpha$, there is no contraction between $v$ and other nodes in $K_{i}\cup\set{\gamma(u_{i})}$, as $\mathcal{D}_{K_{i}\cup\set{\gamma(u_{i})}}$ is non-contracting and we use a perfect merge function to combine this with $\alpha$. Further, for any node $u$ already in the domain of $\alpha$ when $v$ was added, property 3 of perfect merge functions ensures $d_{\alpha}(u,v)\geq\max\set{d_{\alpha}(u,\gamma(u_{i})),d_{\alpha}(v,\gamma(u_{i}))}$. Applying the triangle inequality and appropriate contraction bounds on $\alpha$ prior to this point gives us the result. Notably, Lemma 2 does allow a small amount of contraction, but if metrics in $\mathcal{Y}$ can be “scaled up” and stay in $\mathcal{Y}$ (such as when $\mathcal{Y}$ is the set of HSTs), we immediately get Theorem 3.1 from Lemmas 1 and 2.
Nested embeddings for HSTs
Output: An embedding of $Z_{1}\cup Z_{2}$ into a single $\beta$-HST
1:Let $u=Z_{1}\cap Z_{2}$
2:Create a new HST $T$ that is initially a copy of $T_{1}$
[…]
4:for $i=0$ to the height of $T_{1}$ do
9:      Set $\alpha(x)$ to be the copy of $\alpha_{2}(x)$ created in $T$ at this step
10:Return $\alpha$
We now present a perfect merge function for HSTs. Algorithm 2 takes in two $\beta$-HSTs for a fixed ratio $\beta$ and returns another $\beta$-HST. We show in Lemma 3 that MergeHST is perfect.
Lemma 3
Algorithm 2 is a perfect merge function for embeddings into $\beta$-HSTs.
Figure 1: This figure depicts what happens in the algorithm MergeHST. In particular, $u$ is the node common to both HSTs on the left side of the figure. The triangles represent subtrees, and the letters below the triangles indicate the presence of the corresponding node as a leaf of that subtree. Note that after merging trees, the least common ancestor of $x$ and $y$ is that of $x$ and $u$ from the first tree, as this was higher than that of $y$ and $u$. This is the foundational idea behind why $d_{\alpha}(x,y)\geq\max\set{d_{\alpha_{1}}(u,x),d_{\alpha_{2}}(u,y)}$ holds when $\alpha$ is the output embedding for MergeHST on inputs $\alpha_{1}$ and $\alpha_{2}$.
Figure 1 illustrates how MergeHST works. Given two HSTs that share a common leaf vertex, the function aligns the two HSTs along the path from the shared leaf to the respective roots, and merges the intermediate vertices along the way. By the manner in which we have defined $\beta$-HSTs, the lengths of the edges along the common paths align correctly with each other, enabling the guarantees we require. We defer the correctness proof for this lemma to Appendix 0.A.3. Note that Theorem 3.1 and Lemma 3 plus the fact that all $n$-point metrics probabilistically embed into HSTs with $O(\log n)$ distortion [23] imply Theorem 2.1.
4 Outlier embeddings for HSTs
Now, given the existence of nested embeddings into HSTs, we turn to approximating outlier embeddings into HSTs. In [19], this type of approximation is obtained for $\ell_{2}$ by introducing new “outlier variables” into a semidefinite program for a minimum distortion $\ell_{2}$ embedding and rounding. We will employ a similar technique here using the linear program and associated rounding algorithm of [31] to approximate the optimal distortion for probabilistic HST embeddings.
Let $\Delta$ be the diameter of $(X,d)$ and let $M=\set{2^{0},2,2^{2},\ldots,2^{\lceil\log\Delta\rceil}}$. Let $\zeta$ be the constant factor from Theorem 2.1 and let $B_{i}(r):=\set{j\in X:d(i,j)\leq r}$. We define (HST LP), which is parameterized by inputs $((X,d),k,c)$. The changes from [31] are highlighted in red.
Table 1 gives an intuitive description of the variables in the LP. Note that in our LP $c$ is a parameter, but in [31] it is the variable they are minimizing. The biggest change from [31] is that we have introduced new variables $\delta_{i}$ for each $i\in X$ that we use as an indicator for whether $i$ is an outlier in the embedding.
$\min{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\sum_{i=1}^{n}\delta_{i}}$ (HST LP)
$\forall j,j^{\prime}\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ $\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(4+\zeta\cdot(\delta_{j}+\delta_{j^{\prime}})\cdot\log^{2}k)}\cdot c\cdot d(j,j^{\prime})$ (4)
$\forall i,r,(j,j^{\prime}\in B_{i}(r))\ \ \ \ \ \min(x_{i,j}^{r},x_{ij^{\prime}}^{r})$ $\geq z_{ijj^{\prime}}^{r}$ (5)
$\forall j,r\sum_{i:j\in B_{i}(r)}x_{ij}^{r}$ $=1$ (7)
$\forall j,j^{\prime},(r<d(j,j^{\prime}))\ \ \ \ \ \gamma_{jj^{\prime}}^{r}$ $=1$ (8)
$\forall j,j^{\prime},r\ \ \ \ \gamma_{jj^{\prime}}^{r}$ $\leq 1$ (9)
$\forall i,j,j^{\prime},r\ \ \ \ \ \gamma_{jj^{\prime}}^{r}x_{ij}^{r},z_{ijj^{\prime}}^{r},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta_{i}}$ $\geq 0;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta_{i}\leq 1}$ (10)
variable intuitive meaning
$\delta_{i}$ indicator for whether $i$ is an outlier
$z_{ijj^{\prime}}^{r}$ the probability that $i$ is the representative of $j$ and $j^{\prime}$ at level $r$
$\gamma_{jj^{\prime}}^{r}$ the probability that $j$ and $j^{\prime}$ are separated at level $r$
$x_{ij}^{r}$ the probability that $i$ is the representative of $j$ at level $r$
Table 1: We imagine assigning a “representative” at each node of the HST such that the representative of a leaf node is the leaf node itself, and the representative of any other node is the same as one of its children’s representatives. This table describes the intuitive meaning of each variable in the LP (which is largely the same as in [31]).
Given an optimal solution to this linear program, we need to define an outlier embedding with low distortion and few outliers. To do so, we will first use the rounding algorithm of [31] to find an embedding on the entire metric, and then we take as outliers all vertices $i$ such that $\delta_{i}\geq\delta^{*}$. In Appendix 0.B we formally prove that this rounding algorithm works and gives us Theorem 1.1 as a result. In a similar manner to $\ell_{2}$ outlier embeddings in [19], we use the existence of nested embeddings to bound the optimal LP solution and thus obtain our final result.
Further, if we modify (HST LP) by changing the objective function to $\sum_{i}w_{i}\cdot\delta_{i}$ for some set of weights $w_{i}$, we can use the same technique to obtain a bicriteria approximation for the weighted variant of the outlier embedding problem, where the cost of making $i$ an outlier is $w_{i}$. This result is expressed as Corollary 1, and we prove it in Appendix 0.B.3.